OpenStudy (anonymous):

Find the derivative of y = 12/ sqrt (2x^2 - 9)

OpenStudy (s):

y' = [(0)(2x^2)^(1/2) - (12)(1/2)(2x^2 - 9)^(-1/2) * (4x)] / ((2x^2 - 9)^(1/2))^2 y' = [-(12) * (1/2) * (2x^2 - 9)^(-1/2) * (4x)] / (2x^2 - 9)

OpenStudy (anonymous):

which rule here did you use?? I tried the product rule , which seems easier, but I just got the wrong answer

OpenStudy (s):

Since there is a division you use a quotient rule. [(Derivative of the first term times the second term) minus (the first term times the derivative of the second term)] over (the second term squared)

OpenStudy (s):

First term and second term i mean nominator and denominator

OpenStudy (anonymous):

ohh ok thanks! but is it possible to use product rule if u rearrange the equation as 12 x ( 2x^2 - 9) ^ -1/2

OpenStudy (s):

Hmmm, yea, I think it should work. I would give it a try but I'm on the phone with my girlfriend so I can't do any more math lol =) Sorry. =)

OpenStudy (anonymous):

lol ok thanks

OpenStudy (accessdenied):

Yes, it would work.

OpenStudy (anonymous):

Can you show how to do it in product rule? Cuz I got a wrong answer when i tried..

OpenStudy (accessdenied):

y = 12/ sqrt (2x^2 - 9) y = 12(2x^2 - 9)^(-1/2) y' = 0 (2x^2 - 9)^(-1/2) + 12* (-1/2)(2x^2 - 9)^(-3/2) * (4x) ^ this term disappears = 12 * (-1/2) * (4x) (2x^2 - 9)^(-3/2) = -24x (2x^2 - 9)^(-3/2) = -24x / (2x^2 - 9)^3/2

OpenStudy (s):

I got something like this: y' = (0) * (2x^2 - 9)^(-1/2) + 12 * (-1/2)*(2x^2 - 9)^(-2/3) y' = 12 * (-1/2)*(2x^2 - 9)^(-2/3) y' = (-6) * (2x^2 - 9)^(-2/3)

OpenStudy (s):

Ooo right.. i missed the 4x..

OpenStudy (anonymous):

oh yea, I get it too :) thanks for both of u

OpenStudy (accessdenied):

Yeah, that and it should be to the power -3/2. )Power rule, -1/2 - 1 = -3/2)

OpenStudy (anonymous):

i missed the negative sign haha